# latexFont shape

### v1

@@ 5.2 Quantitative Evaluation
Figure 10: Ablation experiments on the number of referring images.
Application to Existing COD Methods. To verify the generalization of our Ref-COD method, we also apply this idea to existing COD methods. These methods are chosen according to three criteria: a) recently published, b) representative, and c) with open source code. In particular, the methods to participate in the adaptation here include: PFNet [31], PreyNet [24], SINetV2 [1], BSANet [37], BGNet [4], ZoomNet [2], and DGNet [3]. As shown in the third to last row of Tab. I and Fig. 9, the Ref-COD variants of these models achieve better performance than their original versions for all metrics. It is noteworthy that our lightweight baseline model even achieves comparable performance to recent COD methods after incorporating referring images. This indicates that our method does not rely on powerful segmentation networks. The introduction of the reference branch can help significantly improve the performance of a COD model even though it is not strong, reflecting the effectiveness of our Ref-COD idea.
5.3 Ablative Studies
Here, we take our Ref-COD model with image reference, i.e., R2CNet-RefS, as an example to analyze the role of each component in our well-designed framework.
Number of referring images. We also ablate the number of referring images, i.e. $K$, for its impact on the Ref-COD. Considering that there are 5 referring images in each category of the test split of the R2C7K dataset, the value of $K$ changes from 0 to 5. When ‘$50<K<$ 5’, the performance of our R2CNet is calculated by averaging three evaluations, in which this model is guided by $K$ referring images randomly sampled from a certain category of the test split. These experimental results in Fig. 10 show that the performance of our R2CNet improves with the increase of the number of referring images. We believe this is because the acquired common information is less affected by sample differences during this process, and thus a more accurate location of camouflaged targets is achieved.
TABLE II: Ablation experiments on the components of our R2CNet. RMG: the Referring Mask Generation Moudle (Sec. 4.3); RFE: the Referring Feature Enhance Module (Sec. 4.4).
@@ 5.3 Ablative Studies
1 0.772 0.847 0.604 0.044
2 0.800 0.870 0.661 0.038
3 0.792 0.869 0.644 0.040
4 0.805 0.879 0.669 0.036
TABLE III: Ablation experiments on the fusion strategies in the RMG module. ‘DSF’: the fusion between the common representations and visual features; ‘MSF’: the fusion of features at different scales. ‘$\mathcal{F}_{multiply}$’: element-wise multiplication. ‘$\mathcal{F}_{at}(\cdot)$’: affine transformation. ‘$\mathcal{F}_{concate}$’: concatenation. ‘$\mathcal{F}_{clstm}(\cdot)$’: convolutional lstmcell.
@@
TABLE IV: Ablation experiments on the feature enrichment strategy in the RFE module. ‘w/o CSP’ and ‘w/ CSP’: multi-scale enrichment without or with the cross-scale path.
@@
Model Components. We analyze the importance of each component in our R2CNet, and the experimental results are shown in Tab. II. It can be observed that both of our modules can boost the performance of our model, where the RMG module improves the $w$F score from 0.604 to 0.661 and the RFE module improves this score from 0.604 to 0.644. And a more outstanding performance improvement (from 0.604 to 0.669, i.e., 10.8%) appears when we combine them together. These results confirm the effectiveness of the two modules in segmenting camouflaged objects according to the referring images.
Ablation on RMG module. We provide two variants for both the cross-source fusion and the cross-scale fusion of the referring representations and visual features in our RMG module. As shown in Tab. III, our chosen strategy achieves the best performance among the four combinations.
Ablation on RFE module. We investigate four feasible strategies for feature enrichment in our module. These methods include two plug-and-play fashions and two variants of our multi-scale enrichment manner. As can be observed in Tab. IV, our well-designed strategies achieve better performance under reasonable inference speed, especially the one with the cross-scale path.
TABLE V: Ablation experiments on the channel number of our R2CNet.
@@
Model Dimension. The dimension of the model ($c_{d}$) has a significant impact on its parameter scale and inference speed, and we thus conduct several experiments to select a suitable dimension value. Specifically, we increase $c_{d}$ from 32 to 256 in a doubling fashion, and the changes in model parameters, computational cost, and performance during this process are shown in Tab. V. It can be observed that the performance of our R2CNet continues to improve during the increase of $c_{d}$. However, when $c_{d}$ is greater than 64, the improvement of performance begins to decrease, while the parameters and calculation cost of the model increase sharply. For example, when $c_{d}$ increases from 64 to 128, the performance of our R2CNet improves slightly, but the computational overhead increases by 35% (No.2 vs. No.3). Therefore, we set $c_{d}$ to 64 for the trade-off between the performance and the efficiency of our R2CNet.
Ref-Image
Camo-Image
GT
Baseline
Ours–RefS
Ours–RefT
SpiderGhostPipefishLizardPipefishCicadaCatClownfish
Figure 11: Visual comparisons of the predictions between the proposed Ref-COD (R2CNet) and the standard COD (Baseline). ‘-S’: Salient images reference; ‘-T’: Text reference. The segmentation results are shown in purple.
@@ 5.4 Qualitative Evaluation
Results Visualization. We then present the visual comparison of prediction results of our referring COD method, i.e, R2CNet, and the class-agnostic COD method, i.e, Baseline, in Fig. 11. As shown in these scenes from the $1^{st}$ (Spider) to the $4^{th}$ column (Pipefish), there are multiple camouflaged objects. The baseline model tends to indiscriminately segment all potential camouflaged objects, and thus its outputs are often inaccurate. To be specific, it may not only miss some camouflaged objects (e.g., Cicada), but also mispredict other objects as camouflaged ones (e.g., Spider, Lizard). For all these challenging scenes, our R2CNet is able to identify the camouflaged objects orienting to referring images. In addition, compared with the baseline model, our R2CNet can segment visual objects more completely in scenes with only one camouflaged object but with less difference from the surrounding, as shown in the $6^{th}$ column (Cat) and $7^{th}$ column (Clownfish). We guess this is because the common representations of a certain category of objects learned from the referring images are conducive to identifying the main body of such camouflaged objects.
GTw/o Refw/ RefSw/ RefT
Features Visualization. To gain more insight into the Ref-COD, we first visualize intermediate features with or without referring information, as shown in Fig. 12. As can be seen, the model without references can locate the camouflaged objects, nevertheless, it may be disturbed by other objects, and produces inaccurate segmentation results. With the incorporation of references, the model focuses on the indicated camouflaged objects, regardless of other ones.
@@ 6 Future Work
Based on the dataset and framework proposed in this paper, there are a number of foreseeable directions about the Ref-COD to explore for the future research:
1) Interaction between References and Visual Features. Recent works on transformer have proved its great advantages in processing multiple kinds of information. Thus, apart from the direct comparison in this paper, it is promising to explore more effective ways of interaction according to the key component of the transformer architecture, i.e., the attention mechanism.
2) Other Forms of References. In recent years, multimodal research has made great progress in vision and language, vision and speech, etc. Based on them, it is interesting to explore Ref-COD with other references, e.g., text and speech. As discussed above, the main challenge depends on how to obtain these references easily and efficiently.
3) The Extended Tasks. In some application scenarios, users only need some simple answers to make decisions, rather than using more complex equipment to obtain the segmentation details of camouflaged objects. Therefore, it is also valuable to extend the Ref-COD to simple tasks related to question answering.
@@ 7 Conclusions
We propose a novel benchmark named Ref-COD to directionally segment the camouflaged object with simple image references. Fist, we first build a large-scale dataset (i.e., R2C7K) consisting of images in real-world scenarios. Then, we develop an end-to-end Reference to COD framework (i.e., R2CNet), which follows the two-branch architecture. Particularly, the introduction of our well-designed RMG module and RFE module facilitates the guidance of the common representations from referring images to the challenging segmentation of camouflaged objects. Compared with the COD counterpart which indiscriminately segments camouflaged objects from the background, our R2CNet achieves significantly superior performance on the common metrics, and generates more visually favorable predictions. We believe it is promising to establish a multi-information collaborative system, and we hope the perspective provided by our Ref-COD will inspire future work.

### v3

@@ 5.3 Ablative Studies
Figure 10: Ablation studies on the number of referring images. The blue dashed line and orange solid line represent the variation of two COD metrics, i.e., $w$F and $M$, with the number of reference images.
TABLE IV: Ablation studies on the components of our R2CNet. RMG: Referring Mask Generation Moudle. RFE: Referring Feature Enhance Module. Both performance and computational overhead are taken into consideration.
@@
1 0.772 0.847 0.604 0.044 25.97 21.02 185.74
2 0.800 0.870 0.661 0.038 26.40 22.30 169.36
3 0.792 0.869 0.644 0.040 26.55 21.95 166.57
4 0.805 0.879 0.669 0.036 27.15 23.23 151.47
TABLE V: Ablation experiments on the fusion strategies in the RMG module. ‘DSF’: the fusion between the common representations and visual features; ‘MSF’: the fusion of features at different scales. ‘$\mathcal{F}_{multiply}$’: element-wise multiplication. ‘$\mathcal{F}_{at}(\cdot)$’: affine transformation. ‘$\mathcal{F}_{concate}$’: concatenation. ‘$\mathcal{F}_{clstm}(\cdot)$’: convolutional lstmcell.
@@
TABLE VI: Ablation experiments on the feature enrichment strategy in the RFE module. ‘w/o CSP’ or ‘w/ CSP’: multi-scale enrichment without or with the cross-scale path.
@@
Model Components. We analyze the importance of each component in our R2CNet, and the experimental results are shown in Tab. IV. It can be observed that both of our modules can boost the performance of our model, where the RMG module improves the $w$F score from 0.604 to 0.661 and the RFE module improves this score from 0.604 to 0.644. And a more outstanding performance improvement (from 0.604 to 0.669, i.e., 10.8%) appears when we combine them together. These results confirm both of our proposed modules can boost the performance of our model in segmenting camouflaged objects according to the referring images with limited computational overhead.
Ablation on RMG module. We provide two variants for both the cross-source fusion and the cross-scale fusion of the referring representations and visual features in our RMG module. As shown in Tab. V, our chosen strategy achieves the best performance among the four combinations.
Ablation on RFE module. We investigate four feasible strategies for feature enrichment in our module. These methods include two plug-and-play fashions and two variants of our multi-scale enrichment manner. As can be observed in Tab. VI, our well-designed strategies achieve better performance under reasonable inference speed, especially the one with the cross-scale path.
TABLE VII: Ablation experiments on the channel number of our R2CNet.
@@
Model Dimension. The dimension of the model ($c_{d}$) has a significant impact on its parameter scale and inference speed, and we thus conduct several experiments to select a suitable dimension value. Specifically, we increase $c_{d}$ from 32 to 256 in a doubling fashion, and the changes in model parameters, computational cost, and performance during this process are shown in Tab. VII. It can be observed that the performance of our R2CNet continues to improve during the increase of $c_{d}$. However, when $c_{d}$ is greater than 64, the improvement of performance begins to decrease, while the parameters and calculation cost of the model increase sharply. For example, when $c_{d}$ increases from 64 to 128, the performance of our R2CNet improves slightly, but the computational overhead increases by 35% (No.2 vs. No.3). Therefore, we set $c_{d}$ to 64 for the trade-off between the performance and the efficiency of our R2CNet.
[…]
Figure 11: Visual comparisons of the predictions between the proposed Ref-COD method (R2CNet) and the standard COD method (Baseline). The segmentation masks are shown in purple.
TABLE VIII: Ablation experiments on the form of Ref-COD. text-ref: text with a simple prompt; camo-ref: image with camouflaged objects; sal-ref: image with salient objects; LSM: using large-scale pre-trained model; GT: using GT mask of the reference; Seen: the reference representations used in training appears in testing.
[…]
1 Baseline 0.772 0.847 0.604 0.044
2 + text-ref 0.805 0.872 0.661 0.038
3 + camo-ref 0.801 0.869 0.656 0.039
4 + sal-ref 0.805 0.879 0.669 0.036
Reference Forms. We also investigate three different forms of reference information, i.e., text descriptions, images with camouflaged objects, and images with salient objects. As discussed in Sec. 1 and Sec. 2, it is difficult to obtain the detailed text descriptions and the annotated images containing the camouflaged objects associated with the camouflaged targets in the given images. Here we consider the readily available versions of text descriptions and images with camouflaged objects, and compare their upper bounds with our chosen reference, i.e., images with salient objects.
For text reference, inspired by recent prompt engineering, we first build text descriptions with ‘a photo of [CLASS]’, where ‘[CLASS]’ denotes one of the 64 categories in our R2C7K. Then, we feed these 64 text descriptions into the pre-trained CLIP [103] that adopts ResNet-50 as visual backbone, to get the textual features, following CLIPSeg [104]. These features are used as the common representations of target objects. Note that we adopt the same text representations for the samples belonging to the same category in the training set and testing set.
For image reference with camouflaged objects, we first randomly sample several images from the Camo-subset of the category to which the target objects belong. Then, we send them into ResNet-50 to extract visual features. Note that these features are masked and pooled with ground-truth annotations to get the common representations.
The experimental results are shown in Tab. VIII. Note that the text reference employs a large-scale pre-trained model and the common representations in testing have been seen during training. The camouflage candidate adopts GT annotations as masks to capture the common representations. Despite these, our R2CNet using salient object reference still achieves better performance than them. These experiments indicate that referring images with salient objects can more effectively discover and segment the camouflaged objects.
[…]
@@ 5.4 Qualitative Evaluation
Results Visualization. We first present the visual comparison of prediction results of our referring COD method, i.e, R2CNet, and the class-agnostic COD method, i.e, Baseline, in Fig. 11. As shown in these scenes from the $1^{st}$ (Spider) to the $4^{th}$ column (Cicada), there are multiple camouflaged objects. The baseline model tends to indiscriminately segment all potential camouflaged objects, and thus its outputs are often inaccurate. To be specific, it may not only miss some camouflaged objects (e.g., Cicada), but also mispredict other objects as camouflaged ones (e.g., Spider and Lizard). For all these challenging scenes, our R2CNet is able to identify the camouflaged objects orienting to referring images. In addition, compared with the baseline model, our R2CNet is less disturbed by other similar objects (e.g., Pipefish). A nd it can segment visual objects more completely in scenes with only one camouflaged object but with less difference from the surrounding, as shown in the $6^{th}$ column (Cat) and $7^{th}$ column (Clownfish). We argue that this is because the common representations of a certain category of objects learned from the referring images are conducive to identifying the main body of such camouflaged objects.
Features Visualization. To gain more insights from the Ref-COD, we also visualize intermediate features with or without referring information, as shown in Fig. 12. As can be seen, the model without references can locate the camouflaged objects but it may be disturbed by other objects, and produces inaccurate segmentation results. With the incorporation of references, the model can focuse more on the indicated camouflaged objects, regardless of other ones.
@@ 6 Future Work
Based on the R2C7K dataset and R2CNet framework proposed in this paper, there are a few foreseeable directions about the Ref-COD to explore in future research:
1) Other Forms of References. In recent years, multi-modal research has made great progress in vision and language, vision and speech, etc. Based on them, it is interesting to explore Ref-COD with other references, e.g., text and speech. As discussed above, the main challenge depends on how to obtain these references easily and efficiently.
2) Special Case of Referring Scene. In this paper, we assume the image to be segmented contains the target objects specified by the referring images. However, such objects may not exist in some real scenes. It will be promising to extend Ref-COD to be compatible with these cases.
3) Related Tasks. In some application scenarios, users only need some simple answers to make decisions, rather than using more complex equipment to obtain the segmentation details of camouflaged objects. Therefore, it is also valuable to extend the Ref-COD to simple tasks related to question answering.
@@ 7 Conclusions
We propose a novel benchmark named Ref-COD to directionally segment the camouflaged object with simple image references. First, we build a large-scale dataset (i.e., R2C7K) consisting of images in real-world scenarios. Then, we develop an end-to-end Reference to COD framework (i.e., R2CNet), which follows the two-branch architecture. Particularly, the introduction of our well-designed RMG module and RFE module facilitates the guidance of the common representations from referring images to the challenging segmentation of camouflaged objects. Compared with the COD counterpart which indiscriminately segments camouflaged objects from the background, our R2CNet achieves significantly superior performance on the common metrics, and generates more visually favorable predictions. We believe it is promising to establish a multi-information collaborative system, and we hope the perspective provided by our Ref-COD will inspire future work.